The average age of a group of mathematicians and computer scientists is 40. If the mathematicians' average age is 35 and the computer scientists' average age is 50, what is the ratio of the number of mathematicians to the number of computer scientists?
(1) 2.5 (2) 3.5 (3) 2 (4) 3 (5) None of the above

Answer by richwmiller(9135)

mM+cC/(M+C)=40
m=35
c=50
(35M+50C)=40(M+C)
35M+50C=40M+40C
10C=5M
10/5=M/C
2=M/C
(3) 2 answer